A diesel train left the station and travled toward the outer-most station at an average speed of 30 km/h. A cattle train left some time later traveling in the opposite direction with an average speed of 70 km/h. After the trains had traveled for 17 hours they were 790 km apart. How long did the cattle train teavel?
The answer is 4 hours, but I'm not sure how to solve it. Any help is greatly appreciated!

Answer by chana(69)

A diesel train left the station and travled toward the outer-most station at an average speed of 30 km/h. A cattle train left some time later traveling in the opposite direction with an average speed of 70 km/h. After the trains had traveled for 17 hours they were 790 km apart. How long did the cattle train teavel?
here is the solution for your question :
we are going to add up the distance of two trains and equate it to 790 km .
diesel train's speed = 30 km/hr
time travelled = 17 hrs since it started first the total travelled time by both = time taken by train started first
cattle train's speed= 70km/hr
time travelled = (17 - x)hrs----------(1)
dist travelled by first train +dist travelled by cattle train = 790
(30*17) + ((70)*(17-x)) = 790
solving x you get x = 13
sbstitute in 1 you will get required data.